# Waste Heat and Habitability: Constraints from Technological Energy Consumption

### v1

@@ 2.4 Connection to the Kardashev scale
Circling back to the discussion in Section 1, the classic paper by Kardashev (1964) proposed that energy consumption can be used to broadly categorize the technological level of extraterrestrial intelligences. Quantitatively, this classification is expressed using the so-called Kardashev scale, which is mathematically represented as (Gray, 2020):
@@
A frequently adopted prescription for estimating the time required for an advancing technological species to ascend the rungs of the Kardashev scale is to employ a constant exponential growth rate $a$ of energy utilization, [^9] It is known, however, that presupposing any particular growth pattern involves attendant uncertainty (e.g., Smil, 2000), and may not be tenable, as partly elaborated in Section 3. reflecting the observed trend on Earth during the industrial era (von Hoerner, 1975). This assumption can be quantitatively encapsulated as follows:
@@
on simplification, which reveals that $t$ has a modest dependence on $a$ (inverse proportionality) and a very weak (i.e., logarithmic) dependence on $P_{0}$. Even if the latter were to vary by many orders of magnitude, it ought to alter the predicted value of $t$ by a relatively moderate amount (e.g., an order of magnitude), thus conferring some degree of robustness. These two vital points should be borne in mind with respect to upcoming results. For instance, as remarked toward the ending of Section 2.1, reducing the waste heat production – which is crudely tantamount to diminishing the magnitude of $P$ – ought not modify our conclusions substantively.
@@
While the adoption of a constant growth rate may create the impression that a swift transition to ascending levels of the Kardashev scale is bound to happen, our preceding discussion warrants caution. Specifically, we must take into account the inevitable global heating associated with energy consumption (represented by $P$). If, as assumed implicitly in this paper (see Section 2.1), the waste heat flux $Q_{\rm tech}$ broadly obeys the relation,
@@
In the forthcoming Section 2.5, we will further delve into the implications for planetary habitability, and follow this up by calculating the maximal lifetime and Kardashev scale of technological species in Section 2.6.
@@ 2.5 Implications of heating for habitability
We can apply the previous considerations to estimate the ramifications for planetary habitability ensuing from the production of waste heat associated with energy consumption by technological species. In essence, we are interested in forecasting the time it would take for a planet hosting a technological species – pursuing an exponential growth of energy consumption – to experience a rise in temperature so significant to jeopardize its habitability. This matter was studied by the likes of von Hoerner (1975) and Mullan and Haqq-Misra (2019).
In Figure 2, the fractional temperature increase $\Delta T/T_{0}$ is plotted as a function of time lapsed for different annual growth rates of energy assumption by invoking (7), where the heating term is specified appropriately based on the power source (viz., non-stellar vs stellar energy). The two horizontal lines of $\Delta T/T_{0}=0.02$ and $\Delta T/T_{0}=0.15$ represent the crossing of conceivably vital planetary thresholds for habitability justified below. [^12] These thresholds are partially reminiscent of the central “planetary boundaries” and “planetary climate tipping points” concepts in climate science (Lenton et al., 2008; Rockström et al., 2009; Barnosky et al., 2012; Steffen et al., 2015; Raworth, 2017; O’Neill et al., 2018; Ritchie et al., 2021; Persson et al., 2022; Richardson et al., 2023; Rockström et al., 2023; Wunderling et al., 2024), as well as cognate paradigms, but are not equivalent to them. We notice that a major elevation of surface temperature is expected on short timescales of $\lesssim 1000$ years.
@@
However, it is plausible that $\Delta T=12$ K may not represent an accurate upper bound because a potentially crucial factor could be missing, namely, the intricate interplay of ecological interactions. By incorporating such kinds of dependencies into a suite of numerical models and modeling possible trajectories, Strona and Bradshaw (2018) concluded that a swift temperature increase of merely $\Delta T\approx 6$ K might suffice to drive global biodiversity collapse (see also Song et al. 2021; Malanoski et al. 2024 for elaborations of this Earth system picture). State-of-the-art evidence garnered from mass extinctions in the Phanerozoic indicates that a couple of them were characterized by temperature changes of $\sim 10$ K (e.g., Dal Corso et al., 2022; Morgan et al., 2022). Therefore, upon adopting a fiducial temperature of $T_{0}=T_{\oplus}\equiv 288$ K for the planet and heating of $\Delta T\approx 6$ K, we duly obtain $\Delta T/T_{0}\approx 0.02$.
@@ 2.6 Maximum lifespan of technospheres
The circumstellar habitable zone (HZ) is a cornerstone of astrobiology, with roots dating back to the 19th century (Lingam, 2021). The lifetime of the continuous habitable zone – which solely involves abiotic (i.e., neither biological nor technological) factors – was analyzed in several publications (Kasting et al., 1993; Kopparapu et al., 2013; Rushby et al., 2013; Gallet et al., 2017; Levenson, 2021; Ware et al., 2022; Tuchow and Wright, 2023), based on the consideration of a gradually increasing stellar luminosity. We will now proceed to estimate the upper bound on the lifetime of technospheres, where the key determinant of longevity is the energy consumption growth rate of technological species.
In classic models of the HZ, the inner and outer limits are expressed in terms of the corresponding top-of-atmosphere stellar flux $S_{\rm max}$ and $S_{\rm min}$, respectively regulated by the maximum and moist greenhouse effect (Kasting et al., 1993; Kopparapu et al., 2013; Ramirez, 2018; Ramirez et al., 2019; Lingam and Balbi, 2024). These thresholds can, to a certain degree, be perceived as “universal” because they are built from well-defined atmospheric and surface physics and chemistry, and hence they may pose fairly stringent constraints on planetary habitability, although exceptions could exist.
At an initial time (e.g., as when $P=P_{0}$ in Section 2.4), let us suppose that the planet harboring a technological species – albeit one that has not yet entered the exponential growth phase – is at the outer edge of the habitable zone and receives a stellar flux of $S_{\rm min}$. The planet cannot exist any further away from the host star, because liquid water would no longer be viable on the planetary surface (for a given atmospheric composition), and the world may therefore not be capable of sustaining a technological species with an Earth-like biochemistry (viz., entailing liquid water as the solvent).
@@ 3 Discussion and Conclusions
As articulated in Section 1, waste heat production is an inevitable consequence of thermodynamics. In this paper, we investigated the myriad effects of technological species – and their accompanying technospheres – experiencing exponential growth with respect to their energy consumption and waste production.
@@
    Technological species venture beyond their home planet(s), thereby utilizing space infrastructure for producing and dissipating energy, as well as for performing other technological activities.
We shall now tackle each of these trajectories in more detail, and briefly comment on their implications. We emphasize, however, that we do not claim to exhaust all conceivable routes in our précis. A comparatively exhaustive treatment of Earth’s technosphere $1000$ years from the present day, drawing on the domain of future studies, was recently explored in Haqq-Misra et al. (2024).
@@
Trajectory #1: If the majority of technological species are driven to extinction during their pursuit of exponential growth on account of violating one or more planetary thresholds, their mean lifetime during the detectable phase may be markedly low, as indicated earlier. In this specific scenario, the estimated number of communicative technological species in the Galaxy would drop commensurately as per the famous Drake equation formulated by Frank Drake (Drake, 1965), since one of the constituent factors is the average longevity of such species (Mamikunian and Briggs, 1965; Shklovskii and Sagan, 1966; Drake and Sobel, 1992; Burchell, 2006; Shostak, 2011; Vakoch and Dowd, 2015).
If the number of these technological species is indeed modest – based on the Drake equation and its probabilistic descendants (Glade et al., 2012; Maccone, 2012; Mieli et al., 2023), which are regulated by the longevity – this outcome might offer a (partial) “solution” to the famous Fermi paradox (Shermer, 2002; Webb, 2015; Cirkovic, 2018; Forgan, 2019): we have not encountered technological species because they are rare at any given moment in time. This explanation of the Fermi paradox in terms of long-term sustainability – or rather, the lack thereof – was surveyed by the likes of Dutil and Dumas (2009) and Haqq-Misra and Baum (2009), and fleshed out in subsequent publications by Frank et al. (2018); Mullan and Haqq-Misra (2019); Bologna and Aquino (2020); Lingam and Loeb (2021); Savitch et al. (2021); Wong and Bartlett (2022).
@@
Trajectory #2: The pivotal outcomes that could stem from reifying protracted exponential growth of any given resource(s) have attracted intense debates since at least the classic treatise by Malthus (1798). A diverse collection of perspectives advanced by intellectuals spanning multifarious schools of thought – dating back to Mill (1848) at the minimum – have advocated in favor of humankind ultimately shifting to a paradigm of (near-)zero growth of that particular resource(s) sensu lato (i.e., a quasi-steady state in essence) or even embracing appropriate intervals of negative growth, which may be interpreted as degrowth (Georgescu-Roegen, 1971; Daly, 1973; Meadows et al., 1972; Hirsch, 1978; Daly, 1996; Burkett, 2006; Odum, 2007; Latouche, 2009; Kerschner, 2010; Foster et al., 2010; Bardi, 2011; Daly and Farley, 2011; Randers, 2012; D’Alisa et al., 2014; Costanza et al., 2015; Schmelzer, 2015; Schramski et al., 2015; Kallis et al., 2018; Jackson and Victor, 2019; Smil, 2019; Hickel, 2020; Kallis et al., 2020; Millward-Hopkins et al., 2020; Wiedmann et al., 2020; Herrington, 2021; Keyßer and Lenzen, 2021; McPhearson et al., 2021; Foster, 2022; Hickel et al., 2022; Saito, 2022; Schmelzer et al., 2022; Buch-Hansen et al., 2024). It should be recognized, however, that these paradigms have, in turn, garnered a plethora of criticisms, critiques, and responses (e.g., highlighting purported lacunae) from variegated standpoints (Ekins, 2000; Hart, 2005; Martínez-Alier et al., 2010; van den Bergh, 2011, 2017; Anderson and M’Gonigle, 2012; Fay, 2012; Hallegatte et al., 2012; Schwartzman, 2012; Naam, 2013; Quilley, 2013; Cosme et al., 2017; Kish and Quilley, 2017; Obama, 2017; Buch-Hansen, 2018; Liodakis, 2018; Strunz and Bartkowski, 2018; Büchs and Koch, 2019; Spash, 2020; Trainer, 2020; Durand et al., 2024; Gates, 2021; Polewsky et al., 2024; Ritchie, 2024).
For an annual growth rate ($a$) obeying $a\ll 1$, it can be shown that the maximal lifetime of technospheres is proportional to $1/a$ on taking this limit in (21). Hence, if a technological species is capable of drastically reducing the growth rate, its longevity will be proportionally enhanced, although the threshold imposed by (20) is still at play. By the same token, shifting from exponential growth to, say, power-law growth can also extend the lifespan of the technosphere because an exponential function increases faster than a power law in due course; this statement is likewise valid for logistic growth, which is tantamount to having the growth rate approach zero asymptotically. Lastly, the pursuit of growth might engender collapse, but not the extinction, of technological species, thence serving as a negative feedback suppressing growth and elevating their lifespan, albeit with a reduced energy footprint.
@@
To resolve this matter, let us examine (20), which signifies an upper bound on the allowed heat flux. This limit was interpreted as capping the permitted energy consumption (i.e., power) and the Kardashev scale at the end of Section 2.6: refer to (22) and the associated exposition. However, this line of reasoning is implicitly valid only if the total surface area is held fixed. In contrast, if technological species were to expand their activities into space – which has been advocated by many to yield a myriad of benefits to humankind (Bernal, 1929; Stapledon, 1937; Dyson, 1960; O’Neill, 1977; Finney and Jones, 1985; Sagan, 1994; Bostrom, 2003; Harrison, 2001; Dick and Launius, 2007; Beech, 2009; Crawford, 2010; Rovetto, 2013; Munévar, 2014; Friedman, 2015; Crawford, 2016; Shelhamer, 2017; Zubrin, 2019, 2021; Ćirković, 2022; Johnson and Roy, 2023), although this stance has concomitantly garnered its share of critiques (Williams, 2010; Valentine, 2012; Slobodian, 2015; Schwartz and Milligan, 2016; Billings, 2019; Szocik, 2019; Traphagan, 2019; Deudney, 2020; Kovic, 2021; Gunderson et al., 2021; Schwartz et al., 2021) – these entities could decide to deploy space infrastructure for generating energy and dissipating heat over a much wider area (cf. Zotin et al., 2001; Wong, 2017). [^16] From the perspective of creating heat sinks, a distinctly speculative example by Opatrný et al. (2017) proposed that technological species with $K\gtrsim 1$ might control and utilize small black holes for this purpose (see also Loeb 2024). By doing so, the aforementioned bound on the power and the Kardashev scale may be circumvented to some degree.
@@
We caution that unceasing exponential growth, even amidst the much broader horizons of space, might be rendered impractical at some stage in particular cases (as touched on shortly): if so, technological species may eventually adopt a variant of Trajectory #2 and curb their growth. Furthermore, whether Trajectory #3 can be instantiated while conceivably grappling with the specter of Trajectory #1 at a certain juncture (e.g., early on in history) could constitute an open question. An interesting feature of partial relevance in this context is the timescales corresponding to various phenomena. We have shown in this paper that the exponential growth of putative technological species on Earth-like planets can become unsustainable in $\sim 10^{3}$ yr under specific assumptions; whereas the duration for effectuating self-sustaining and stable planetary-scale modification (terraforming) of another world in toto might be up to $\lesssim 10^{5}$ yr with current human capabilities (cf. McKay and Marinova, 2001; Beech, 2009), for reaching neighboring stars with chemical propulsion is conventionally $\sim 10^{5}$ yr, and for settling the entire Milky Way is estimated to be $\sim 10^{7}$$10^{9}$ yr (Jones, 1976, 1981; Cotta and Morales, 2009; Nicholson and Forgan, 2013; Lingam, 2016; Carroll-Nellenback et al., 2019; Haqq-Misra and Fauchez, 2022).
@@
Coda: To round off this discourse, two of the three core themes underpinning astrobiology (arguably comprising its raison d’être) are centered on the fundamental questions of: (1) the putative existence and distribution of extraterrestrial life; and (2) the future of life on Earth and beyond (e.g., Chyba and Hand, 2005; Des Marais et al., 2008; Domagal-Goldman et al., 2016; Lingam and Balbi, 2024). By employing simple thermodynamic models, we have attempted to gauge the maximal lifetime of putative technological species and their technospheres characterized by exponential growth of energy consumption (a category which crucially encompasses humans in this epoch), since this endeavor has a direct bearing on the preceding two themes in this discipline.
Hence, even while several avenues that evidently warrant further research beyond this preliminary undertaking are apparent – such as notably harnessing sophisticated climate models to self-consistently calculate the temporal climate (co)evolution of temperate terrestrial planets hosting technospheres fuelled by sundry sources of energy or incorporating alternative modes of energy growth (e.g., logistic or power law growth) – which might consequently refine or modify (or perhaps even contradict) our central findings, we nevertheless hope that this work may serve to capture some of the salient physical details, and pave the way for in-depth treatments to chart the trajectories accessible by technological species in the Universe, including humans, as well as the ensuing signatures of technological activities (viz., technosignatures).

### v2

@@ 2.4 Connection to the Kardashev scale
Circling back to the discussion in Section 1, in his classic paper, Kardashev (1964) proposed that energy consumption can be used to broadly categorize the technological level of extraterrestrial intelligences. Quantitatively, this classification is expressed using the so-called Kardashev scale, which is mathematically represented as (Gray, 2020):
@@
A frequently adopted prescription for estimating the time required for an advancing technological species to ascend the rungs of the Kardashev scale is to employ a constant exponential growth rate $a$ of energy utilization, [^9] It is known, however, that presupposing any particular growth pattern involves attendant uncertainty (e.g., Smil, 2000) and may not be tenable, as partly elaborated in Section 3. reflecting the observed trend on Earth during the industrial era (von Hoerner, 1975). This assumption can be quantitatively encapsulated as follows:
@@
on simplification, which reveals that $t$ has a modest dependence on $a$ (inverse proportionality) and a very weak (i.e., logarithmic) dependence on $P_{0}$. Even if the latter were to vary by many orders of magnitude, it ought to alter the predicted value of $t$ by a relatively moderate amount (e.g., an order of magnitude), thus conferring some degree of robustness. These two vital points should be borne in mind with respect to upcoming results. For instance, as remarked toward the end of Section 2.1, reducing the waste heat production – which is crudely tantamount to diminishing the magnitude of $P$ – ought not modify our core conclusions substantively.
@@
While the adoption of a constant growth rate may create the impression that a swift transition to ascending levels of the Kardashev scale is bound to happen, our preceding discussion warrants caution. Specifically, we must take into account the inevitable global heating associated with energy consumption (represented by $P$). If, as assumed implicitly in this paper (see Section 2.1), the waste heat flux $Q_{\rm tech}$ broadly obeys the relation
@@
In the forthcoming Section 2.5, we will further delve into the implications for planetary habitability, and we will follow this up by calculating the maximal lifetime and Kardashev scale of technological species in Section 2.6.
@@ 2.5 Implications of heating for habitability
We can apply the previous considerations to estimate the ramifications for planetary habitability that ensue from the production of waste heat associated with energy consumption by technological species. In essence, we are interested in forecasting the time it would take for a planet that hosts a technological species – which pursues an exponential growth of energy consumption – to experience a rise in temperature so significant as to jeopardize its habitability. This matter was studied by the likes of von Hoerner (1975) and Mullan and Haqq-Misra (2019).
In Figure 2, the fractional temperature increase $\Delta T/T_{0}$ is plotted as a function of time lapsed (since the commencement of exponential growth) for different annual growth rates of energy assumption by invoking (7), where the heating term is specified appropriately based on the power source (viz., non-stellar vs stellar energy). The two horizontal lines of $\Delta T/T_{0}=0.02$ and $\Delta T/T_{0}=0.15$ represent the crossing of conceivably vital “planetary thresholds” for habitability justified below. [^12] These thresholds are partially reminiscent of the central “planetary boundaries” and “planetary climate tipping points” concepts in climate science (Lenton et al., 2008; Rockström et al., 2009; Barnosky et al., 2012; Steffen et al., 2015; Raworth, 2017; O’Neill et al., 2018; Ritchie et al., 2021; Persson et al., 2022; Richardson et al., 2023; Rockström et al., 2023, 2024; Wunderling et al., 2024), as well as cognate paradigms, but are not equivalent to them. We notice that a major elevation of surface temperature is expected on short timescales of $\lesssim 1000$ years.
@@
However, it is plausible that $\Delta T=12$ K may not represent an accurate upper bound because a potentially crucial factor could be missing, namely, the intricate interplay of ecological interactions. By incorporating such kinds of dependencies into a suite of numerical models and modeling possible trajectories, Strona and Bradshaw (2018) concluded that a swift temperature increase of merely $\Delta T\approx 6$ K might suffice to drive global biodiversity collapse (see also Song et al. 2021; Malanoski et al. 2024 for elaborations of this Earth system picture). State-of-the-art evidence garnered from mass extinctions in the Phanerozoic indicates that a couple of them were characterized by temperature changes of $\sim 10$ K (e.g., Dal Corso et al., 2022; Morgan et al., 2022). Therefore, when we adopt a fiducial temperature of $T_{0}=T_{\oplus}\equiv 288$ K for the planet and heating of $\Delta T\approx 6$ K, we duly obtain $\Delta T/T_{0}\approx 0.02$.
@@ 2.6 Maximum lifespan of technospheres
The circumstellar habitable zone (HZ) is a cornerstone of astrobiology; its roots date back to the 19th century (Lingam, 2021). The lifetime of the continuous habitable zone – which solely involves abiotic (i.e., neither biological nor technological) factors – was analyzed in several publications (Kasting et al., 1993; Kopparapu et al., 2013; Rushby et al., 2013; Gallet et al., 2017; Levenson, 2021; Ware et al., 2022; Tuchow and Wright, 2023), based on the consideration of a gradually increasing stellar luminosity. We will now proceed to estimate the upper bound on the lifetime of technospheres, where the key determinant of longevity is the energy consumption growth rate of technological species.
In classic models of the HZ, the inner and outer limits are expressed in terms of the corresponding top-of-atmosphere stellar flux $S_{\rm max}$ and $S_{\rm min}$, respectively regulated by the maximum and moist greenhouse effect (Kasting et al., 1993; Kopparapu et al., 2013; Ramirez, 2018; Ramirez et al., 2019; Lingam and Balbi, 2024). These thresholds can, to a certain degree, be perceived as “universal” because they are built from well-defined atmospheric and surface physics and chemistry; hence they may pose fairly stringent constraints on planetary habitability, although exceptions could exist.
At an initial time (e.g., as when $P=P_{0}$ in Section 2.4), let us suppose that the planet harboring a technological species – albeit one that has not yet entered the exponential growth phase – is at the outer edge of the habitable zone and receives a stellar flux of $S_{\rm min}$. The planet cannot exist any further away from the host star, because liquid water would no longer be viable on the planetary surface (for a given atmospheric composition), and the world may therefore not be capable of sustaining a technological species with an Earth-like biochemistry (viz., with liquid water as the solvent).
@@ 3 Discussion and Conclusions
As articulated in Section 1, waste heat production is an inevitable consequence of thermodynamics. In this paper, we investigated the myriad effects of technological species – and their accompanying technospheres – that experience exponential growth with respect to their energy consumption and waste production.
@@
    Technological species venture beyond their home planet(s), thereby utilizing space infrastructure to produce and dissipate energy, as well as to perform other technological activities.
We shall now tackle each of these trajectories in more detail and briefly comment on their implications. We emphasize, however, that we do not claim to exhaust all conceivable routes in our précis. A comparatively exhaustive treatment of Earth’s technosphere $1000$ years from the present day, which draws on the domain of future studies, was recently explored in Haqq-Misra et al. (2024).
@@
Trajectory #1: In case the majority of technological species are driven to extinction over the course of their postulated interval of exponential growth on account of violating one or more “planetary thresholds” (sketched in Section 2.5), their mean lifetime during the detectable phase might become markedly low, as intimated earlier. In this specific scenario, the estimated number of communicative technological species in the Galaxy would drop commensurately as per the famous Drake equation formulated by Frank Drake (1965), since one of the constituent factors is the average longevity of such species (Mamikunian and Briggs, 1965; Shklovskii and Sagan, 1966; Drake and Sobel, 1992; Burchell, 2006; Shostak, 2011; Vakoch and Dowd, 2015).
If the number of these technological species is indeed modest – based on the Drake equation and its probabilistic descendants (Glade et al., 2012; Maccone, 2012; Mieli et al., 2023), which are regulated by the longevity – this outcome might offer a (partial) “solution” to the famous Fermi paradox (Shermer, 2002; Webb, 2015; Cirkovic, 2018; Forgan, 2019): we have not encountered technological species because they are rare at any given moment in time. This explanation of the Fermi paradox in terms of long-term sustainability – or rather, the lack thereof – was surveyed by the likes of Dutil and Dumas (2009) and Haqq-Misra and Baum (2009) and fleshed out in subsequent publications by Frank et al. (2018); Mullan and Haqq-Misra (2019); Bologna and Aquino (2020); Lingam and Loeb (2021); Savitch et al. (2021); Wong and Bartlett (2022).
@@
Trajectory #2: The pivotal outcomes that could stem from reifying protracted exponential growth of any given resource(s) have attracted intense debates since at least the classic treatise by Malthus (1798). A diverse collection of perspectives advanced by intellectuals spanning multifarious schools of thought – dating back to Mill (1848) at the minimum – has advocated in favor of humankind ultimately shifting to a paradigm of (near-)zero growth of that particular resource(s) sensu lato (i.e., a quasi-steady state in essence) or even embracing suitable intervals of negative growth, which may be interpreted as degrowth (Georgescu-Roegen, 1971; Daly, 1973; Meadows et al., 1972; Hirsch, 1978; Daly, 1996; Burkett, 2006; Odum, 2007; Latouche, 2009; Kerschner, 2010; Foster et al., 2010; Bardi, 2011; Daly and Farley, 2011; Randers, 2012; D’Alisa et al., 2014; Costanza et al., 2015; Schmelzer, 2015; Schramski et al., 2015; Kallis et al., 2018; Jackson and Victor, 2019; Smil, 2019; Hickel, 2020; Kallis et al., 2020; Millward-Hopkins et al., 2020; Wiedmann et al., 2020; Herrington, 2021; Keyßer and Lenzen, 2021; McPhearson et al., 2021; Foster, 2022; Hickel et al., 2022; Saito, 2022; Schmelzer et al., 2022; Buch-Hansen et al., 2024). It should be recognized, however, that these paradigms have garnered a plethora of criticisms, critiques, and responses (e.g., highlighting purported lacunae) from multiple standpoints (Ekins, 2000; Hart, 2005; Martínez-Alier et al., 2010; van den Bergh, 2011, 2017; Anderson and M’Gonigle, 2012; Fay, 2012; Hallegatte et al., 2012; Schwartzman, 2012; Naam, 2013; Quilley, 2013; Cosme et al., 2017; Kish and Quilley, 2017; Obama, 2017; Buch-Hansen, 2018; Liodakis, 2018; Strunz and Bartkowski, 2018; Büchs and Koch, 2019; Spash, 2020; Trainer, 2020; Durand et al., 2024; Gates, 2021; Polewsky et al., 2024; Ritchie, 2024).
For an annual growth rate ($a$) obeying $a\ll 1$, it can be shown that the maximal lifetime of technospheres is proportional to $1/a$ on taking this limit in (21). Hence, if a technological species is capable of drastically reducing the growth rate, its longevity will be proportionally enhanced, although the threshold imposed by (20) is still at play. By the same token, shifting from exponential growth to, say, power-law growth can also extend the lifespan of the technosphere because an exponential function increases faster than a power law in due course; this statement is likewise valid for logistic growth, which is tantamount to having the growth rate approach zero asymptotically. Lastly, the pursuit of growth might engender collapse, but not extinction, of technological species, thence serving as a negative feedback that suppresses growth and elevates their lifespan, albeit with a reduced energy footprint.
@@
To resolve this matter, let us examine (20), which signifies a rough upper bound on the allowed heat flux. This limit was interpreted as capping the permitted energy consumption (i.e., power) and the Kardashev scale at the end of Section 2.6: refer to (22) and the associated exposition. However, this line of reasoning is implicitly valid only if the total surface area is held fixed. In contrast, if technological species were to expand their activities into space – which has been advocated by many to yield a myriad of benefits to humankind (Bernal, 1929; Stapledon, 1937; Dyson, 1960; O’Neill, 1977; Finney and Jones, 1985; Sagan, 1994; Bostrom, 2003; Harrison, 2001; Dick and Launius, 2007; Beech, 2009; Crawford, 2010; Rovetto, 2013; Munévar, 2014; Friedman, 2015; Crawford, 2016; Shelhamer, 2017; Zubrin, 2019, 2021; Ćirković, 2022; Johnson and Roy, 2023), although this stance has concomitantly garnered its share of critiques (Williams, 2010; Valentine, 2012; Slobodian, 2015; Schwartz and Milligan, 2016; Billings, 2019; Szocik, 2019; Traphagan, 2019; Deudney, 2020; Kovic, 2021; Gunderson et al., 2021; Schwartz et al., 2021) – these entities could decide to deploy space infrastructure to generate energy and dissipate heat over a much wider area (cf. Zotin et al., 2001; Wong, 2017). [^16] From the perspective of creating heat sinks, an extremely speculative example by Opatrný et al. (2017) proposed that technological species with $K\gtrsim 1$ might control and utilize small black holes for this purpose (see also Loeb 2024).
By doing so, the aforementioned bound(s) on the power and the Kardashev scale may be circumvented to some degree. A patently similar scenario involves transferring various energy-intensive processes and activities off-world, not to empty space as such, but to uninhabited and uninhabitable objects instead: on these worlds – which can constitute a subset of the putative so-called “service worlds” (Wright et al., 2022) – exceeding the prior flux and power thresholds may not pose significant impediments to implementing certain energetic components of a technosphere (e.g., stellar energy transduction) owing to the posited absence of habitable conditions and/or a biosphere to begin with.
@@
We caution that unceasing exponential growth, even amidst the much broader horizons of space, might be rendered impractical at some stage in particular cases (as touched on shortly): if so, technological species may eventually adopt a variant of Trajectory #2 and curb their growth. Furthermore, whether Trajectory #3 can be instantiated while conceivably grappling with the specter of Trajectory #1 at a certain juncture (e.g., early on in history) could constitute an open question. An interesting feature of partial relevance in this context is the timescales corresponding to various phenomena. We have shown in this paper that the exponential growth of putative technological species on Earth-like planets can become unsustainable in $\sim 10^{3}$ yr under specific assumptions, whereas the duration for effectuating self-sustaining and stable planetary-scale modification (terraforming) of another world in toto might be up to $\lesssim 10^{5}$ yr with current human capabilities (cf. McKay and Marinova, 2001; Beech, 2009), for reaching neighboring stars with chemical propulsion is conventionally $\sim 10^{5}$ yr, and for settling the entire Milky Way is estimated to be $\sim 10^{7}$$10^{9}$ yr (Jones, 1976, 1981; Cotta and Morales, 2009; Nicholson and Forgan, 2013; Lingam, 2016; Carroll-Nellenback et al., 2019; Haqq-Misra and Fauchez, 2022).
@@
Coda: To round off this discourse, two of the three core themes that underpin astrobiology (arguably comprising its raison d’être) are centered on the fundamental questions of (1) the potential existence and distribution of extraterrestrial life and (2) the future of life on Earth and beyond (e.g., Chyba and Hand, 2005; Des Marais et al., 2008; Domagal-Goldman et al., 2016; Lingam and Balbi, 2024). By employing simple thermodynamic models, we have attempted to gauge the maximal lifetime of putative technological species and their technospheres characterized by exponential growth of energy consumption (a category which crucially encompasses humans in this epoch), since this endeavor has a direct bearing on the preceding two themes in this discipline.
Hence, even while several avenues that evidently warrant further research beyond this preliminary undertaking are apparent – such as notably harnessing sophisticated climate models to self-consistently calculate the temporal climate (co)evolution of temperate terrestrial planets that host technospheres fuelled by sundry sources of energy or incorporating alternative modes of energy growth (e.g., logistic or power law growth) – which might consequently refine or modify (or perhaps even contradict) our central findings, we nevertheless hope that this work may serve to capture some of the salient physical details and pave the way for in-depth treatments to chart the trajectories accessible by technological species in the universe, including humans, as well as the ensuing signatures of technological activities (viz., technosignatures).